Assuming an interest rate of 4.7 percent, what is the value of the following cash flows five years from today?

Year 1: 3715

Year 2: 4795

Year 3: 5805

Year 4: 7120
Answer #1

Solution-

A= P(1+r) ^n

A= Future value

P= Present value

r= Rate of interest

n= number of years

A= 3715*(1.047)^4+ 4795*(1.047)^3 + 5805*(1.047)^2 + 7120*(1.047)^1

A= 4464.2195 + 5503.3692 + 6363.4932 + 7454.64= $23,785.7219

future value of cash flows = $23,785.72 (approx)